The sum of two numbers is 29. Their difference is 7. What are the two numbers?
Answer by checkley71(8403)

x+x+7=29
2x=29-7
2x=22
x=22/2
x=11 answer for the smaller number.
11+7=18 for the larger number.
proof
11+18=29
29=29